OpenStudy (zzr0ck3r):

If \(T: \mathbb{R}^3\rightarrow \mathbb{R}^3\) is a linear transformation of rank \(2\) , then rank \(T^2 \le 1\) . Give proof if true, else give a counter example.

OpenStudy (zzr0ck3r):

Looking over old masters exams and I am not sure on this...

ganeshie8 (ganeshie8):

I think we may choose T such that T^2 = T ?

OpenStudy (kainui):

Yeah I was gonna say the same thing, specifically just pick some vector and use it as the projection operator since that's easy to visualize as going from 3D space to a 2D surface and then doing the same mapping is essentially the identity operation now.

OpenStudy (kainui):

Oh I should clarify since using a single vector as a projection operation in 3D space is rank 1 and you want rank 2. Luckily every vector defines a unique 2D surface in 3D space so: Let's say the projecion operator is \(P = \frac{vv^\top}{v^\top v}\) quick to show \(P^2=P\). And easy to pick some random vector v to make it work out. Then you just need to produce a counter example so no rigorous proof is necessary once you have the thing, just gotta show that \(T^2\) is not rank 1 or less after showing T is rank 2. So as I was talking about with the vector defining a plane: \[T = I-P\] \[T^2 = (I-P)^2 = I^2 -2P + P^2 = I- 2P+P = I-P=T\] So \(T^2\) and \(T\) have the same rank.

ganeshie8 (ganeshie8):

Nice nice. so all projection matrices which project a vector onto a 2D plane are counterexamples.

OpenStudy (bobo-i-bo):

An explicit example is a linear map with this corresponding matrix: \[ \left( \begin{array}{ccc} 1 & 0 & 0 \\ 0 & 1 & 0 \\ 0 & 0 & 0 \end{array} \right)\]

OpenStudy (zzr0ck3r):

ha, sorry. After I posted this I could not get back online.